Two years ago you acquired a 10 year zero coupon bond,$1000 par value bond at 12% YTM.Recently you sood this bond at 8% YTM. Using a semi-annual compounding,compute the annualized horizon for this investment.
1
Expert's answer
2021-09-15T11:27:20-0400

Solution:

Annualized horizon return is the total annual return received on a bond made over a horizon period. It measures the rate of return if the bond is being sold prior to its maturity. In terms of a zero-coupon bond, the horizon return will be calculated as follows:


Interest = "\frac{12\%\;annually}{2\;interest \;payments\;per\;year} =6\%"


Semi-annual periods = 2 interest payments per year "\times" 10 years = 20 semi-annual periods

Purchase price = Face value "\times" PVF of 6"\%" = 1000 "\times" 0.31180 = 311.80

 

After 2 years, the selling price will be calculated as follows:

Interest = "\frac{8\%\;annually}{2\;interest \;payments\;per\;year} =4\%"


Semi-annual periods = 2 interest payments per year "\times" 8 years = 16 semi-annual periods

Selling price = Face value"\times"PVF of 4"\%" = 1000 "\times"0.53391 = 533.91